OpenStudy (anonymous):

calculate the average rate of change for the function f(x)=-x^4+4x^3-2x^2+x+1 from x=0 to x=1

OpenStudy (zarkon):

compute\[\frac{f(1)-f(0)}{1-0}\]

OpenStudy (anonymous):

so the answer i 1? @Zarkon

OpenStudy (agent0smith):

No, show your work.

OpenStudy (anonymous):

f(1)-f(0)/1-0 that would be 0 @agent0smith

OpenStudy (anonymous):

MERRY CHRISTMAS EVERYBODY!!

OpenStudy (agent0smith):

No. When I said show your work, I mean, find f(1), and find f(0). What is f(1) = ? What is f(0) = ?

OpenStudy (anonymous):

f(1)=2 f(0)=0 @agent0smith

OpenStudy (agent0smith):

You're using -x^4+4x^3-2x^2+x+1 and plugging in x=1 for f(1) right? Those aren't correct.

OpenStudy (anonymous):

oh no i wasnt!! well i just sis it now and got 7? that seems weird though so i hope its right

OpenStudy (agent0smith):

Then... how did you get f(1) and f(0)..??? 7 isn't right. What is f(1) = ? What is f(0) = ?

OpenStudy (anonymous):

i thought 1 and 0

OpenStudy (agent0smith):

f(1) means plug in x=1 into the function, -x^4+4x^3-2x^2+x+1. f(0) means plug in x=0 into the function, -x^4+4x^3-2x^2+x+1.

OpenStudy (anonymous):

ohh so its both 1!!

OpenStudy (agent0smith):

er.. what's both?

OpenStudy (anonymous):

i mean they both equal 1

OpenStudy (agent0smith):

No. \[\Large f(1)=-(1^4)+4(1)^3-2(1)^2+1+1 = ???\]

OpenStudy (agent0smith):

-1+4-2+1+1 = 3

OpenStudy (anonymous):

what do we do next

OpenStudy (agent0smith):

Now from that, can you tell me what f(1) is? And f(0)?

OpenStudy (anonymous):

1 and 2

OpenStudy (agent0smith):

No, f(1) is not 1, f(0) is not 2. I already calculated f(1) above for you.

OpenStudy (agent0smith):

Show EXACTLY what you are doing to find f(0) and f(1). Like I did above.

OpenStudy (agent0smith):

\[\Large f(1)=-(1^4)+4(1)^3-2(1)^2+1+1 = ???\] \[\Large f(0)=-(0^4)+4(0)^3-2(0)^2+0+1 = ???\]

OpenStudy (charlotte123):

@agent0smith I wish you luck - You are going to need It...........badly >.>

OpenStudy (anonymous):

that wasnt nice

OpenStudy (charlotte123):

@helpineedit Everyone needs luck! :)